Find the equation of the right circular cone vwhen the straight line 2y+3z=6, x=0 revolves about the z-axis
1
Expert's answer
2020-11-23T13:40:37-0500

Given the vertex point "p_0=(x_0,y_0,z_0)=(0,0,2)," and "p=(x,y,z)"


the support circle in the "xy" plane is "x^2+y^2=r^2" with "r=3."


The generatrix line is given as "\frac{x-x_0}{l}=\frac{y-y_0}{m}=\frac{z-z_0}{n}" so its intersection with the "xy"  plane is given by

"(x_0-z_0\frac{l}{n},y_0-z_0\frac{m}{n},0)"


so this point will lie on the given circle "(x_0-z_0\frac{l}{n})^2+(y_0-z_0\frac{m}{n})^2=r^2".


Here we have "(2\frac{l}{n})^2+(2\frac{m}{n})^2=9"


but "\begin{cases}l=\frac{x}{k}\\m=\frac{y}{k}\\n=\frac{z-2}{k} \end{cases}" substituting


"(2\frac{x}{z-2})^2+(2\frac{y}{z-2})^2=9"


and finally


"\frac{x^2}{9}+\frac{y^2}{9}-\frac{(z-2)^2}{4}=0".